Square tiles measuring 2 feet on each side are sold in bundles of 10 for $13 what will it cost to cover a rectangle floor measuring 40 feet by 50 feet with the square tiles copuld you show me how to set this up and solve it?

Answer by checkley79(3341)

ONE SQUARE IS 2*2=4 SQUARE FEET.
40*50=2,000 SQUARE FEET.
2000/4=500 TILES.
500/10=50 BOXES OF TILE ARE REQUIRES.
50*13=$650.00 FOR THE TILE.